# New Bounds on Quotient Polynomials with Applications to Exact Divisibility and Divisibility Testing of Sparse Polynomials

### v3

@@ 3 A Bound on the $\ell_{2}$-norm of the Quotient Polynomial
Let $g\in\mathbb{C}[x]$ be a monic polynomial. Then, there exists a set $B(g)\subseteq[0,2\pi)$ of points, of size less than $12(\left\|{g}\right\|_{0}-1)\deg{g}$, such that for every $\delta>0$ and every $\alpha\in[0,2\pi)$ that is $(\left\|{g}\right\|_{0}-1)\delta$-far from $B(g)$ it holds that $\left|g(e^{i\alpha})\right|\geq\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}$.
@@
Observe that $\left|(g/x^{\text{ord}_{0}{g}})(\omega)\right|=\left|g(\omega)\right|$ for every $\omega$ on the unit circle and so we can assume without loss of generality that $\text{ord}_{0}{g}=0$. We prove this lemma by induction on $s=\left\|{g}\right\|_{0}$. The claim holds for $s=1$ since it implies that $g=1$, and so clearly $|g(e^{i\alpha})|=1$ for any $\alpha$. For the inductive step, suppose that the claim holds for $s$, and let $g\in\mathbb{C}[x]$ be of sparsity $s+1$. Let $f:[0,2\pi)\to\mathbb{C}$ defined as $f(x)=g(e^{ix})$. Let,
@@
$B(g)$ $=B(g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}}))\cup\left\{\alpha\in[0,2\pi):h(\alpha)=0\text{ for a non-zero }h\in H\right\}.$
By 2.6, each non-zero function in $H$ has at most $2\deg{g}$ zeros in $[0,2\pi)$. Since $g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}})$ is clearly of sparsity $s$, by the inductive assumption we conclude that
$\left|B(g)\right|<12(s-1)\deg{g^{\prime}}+12\deg{g}<12s\deg{g}.$
Let $I=[a,b]\subseteq[0,2\pi)$ be an interval that is $(s-1)\delta$-far from $B(g)$. By definition of $B(g)$, each of the functions in the set $H$ defined in Equation (3.8) does not change its sign in $I$ (as every sign change would have to go through a zero of the function). As $f^{\prime}(x)=ie^{ix}\cdot{g^{\prime}}(e^{ix})$, 3.6 implies that for every $\beta\in[a+\delta,b-\delta]$ it holds that
$\left|f(\beta)\right|$ $\geq\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|f^{\prime}(\alpha)\right|=\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|g^{\prime}(e^{i\alpha})\right|=\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|(g^{\prime}/x^{\text{ord}_{0}{g^{\prime}}})(e^{i\alpha})\right|$
$\geq\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|(g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}}))(e^{i\alpha})\right|.$
As $B(g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}}))\subseteq B(g)$ and each $\alpha\in I$ is at least $(s-1)\delta$-far from $B(g)$, we get from the induction hypothesis applied to $g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}})$ that $\left|f(\beta)\right|\geq\left(\frac{\delta}{\sqrt{2}}\right)^{s}$. In other words, every $\beta$ that is $s\delta$ far from the set $B(g)$ satisfies $\left|f(\beta)\right|\geq\left(\frac{\delta}{\sqrt{2}}\right)^{s}$, which proves the step of the induction. ∎
@@
Let $g\in\mathbb{C}[x]$ be a monic polynomial. Let $p_{min}\in\mathbb{N}$ and $L(n)=n\log{n}$. Set $p_{max}=\lceil 2L(p_{min}+12\left\|{g}\right\|_{0}\deg{g})\rceil$. Then, there exists a prime $p\in(p_{min},p_{max}]$ such that
$\left|g(\omega)\right|\geq\left(\frac{\pi}{\sqrt{2}\left\|{g}\right\|_{0}p_{max}^{2}}\right)^{\left\|{g}\right\|_{0}-1}\geq\left(\frac{2}{\left\|{g}\right\|_{0}p_{max}^{2}}\right)^{\left\|{g}\right\|_{0}-1},$
@@
If $\left\|{g}\right\|_{0}=1$, then the statement is trivial since $\left|g(\omega)\right|=1$ for every root of unity $\omega$. Hence, we can assume that $\left\|{g}\right\|_{0}\geq 2$ and $\deg{g}\geq 1$.
Let $t=p_{min}+12\left\|{g}\right\|_{0}\deg{g}$, and let $\pi(n)$ be the prime counting function. Our assumption implies that $t\geq 24$. By 2.8 we conclude that
$\pi(p_{max})$ $=\pi(\lceil 2t\log{t}\rceil)$
@@
Let $B(g)$ be as in 3.7 and
$P=\left\{p_{min}<p\leq p_{max}:p\text{ is prime}\right\}.$
We bound the size of $P$ by counting the primes in the interval $(p_{min},p_{max}]$:
$\left|P\right|\geq\pi(p_{max})-p_{min}\geq t-p_{min}=12\left\|{g}\right\|_{0}\deg{g}>\left|B(g)\right|.$
By 3.2, there exists $p\in P$ such that for any integer $0<a<p$ and $\alpha\in B(g)$
$\left|a/p-\alpha/2\pi\right|\geq\frac{1}{2p_{max}^{2}}.$
@@
$\left|2\pi\frac{a}{p}-\alpha\right|\geq\frac{\pi}{p_{max}^{2}}.$
Finally, by applying 3.7 on $\delta=\frac{\pi}{\|g\|_{0}p_{max}^{2}}$, we conclude that for any $p$th root of unity $\omega\neq 1$,
$\left|g(\omega)\right|\geq\left(\frac{\pi}{\sqrt{2}\left\|{g}\right\|_{0}p_{max}^{2}}\right)^{\left\|{g}\right\|_{0}-1}.\qed$
@@
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$. Denote $T(n)=n\log{n}$. Then,
$\left\|{h}\right\|_{2}\leq\sqrt{2}\frac{\left\|{f}\right\|_{1}}{LC(g)}\left(3\left\|{g}\right\|_{0}\cdot L\left(12\left\|{g}\right\|_{0}\deg{g}+2\deg{f}\right)^{2}\right)^{\left\|{g}\right\|_{0}-1}.$
@@
First, we can assume that $g$ is monic: if $g$ is not monic, then we scale both $f$ and $g$ by $\frac{1}{LC(g)}$. Let $p_{min}=2\deg{f}$ and $p_{max}=\lceil 2L(12\|g\|_{0}\deg{g}+2\deg{f})\rceil$. By plugging those values into 3.9 we conclude that there exists a prime $p\in(p_{min},p_{max}]$ such that
$\left|g(\omega)\right|\geq\left(\frac{2}{\left\|{g}\right\|_{0}\cdot(\lceil 2L(12\left\|{g}\right\|_{0}\deg{g}+2\deg{f})\rceil)^{2}}\right)^{\left\|{g}\right\|_{0}-1}$
@@
Let $d_{1},d_{2}\in\mathbb{N}$ and $a\in\mathbb{C}$ with $\left|a\right|\geq 1$. Then,
$x^{d_{1}}\bmod x^{d_{2}+1}-ax^{d_{2}}=a^{d_{1}-d_{2}}x^{d_{2}}\qed$
The proof of Theorem 1.10 follows immediately from the bound proved in Theorem 3.10 and the analysis of the long division algorithm given in Section 5.
4 A Bound on the Sparsity of the Quotient Polynomial
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. Throughout this section we assume, without loss of generality, that $\text{ord}_{0}{g}=0$. Indeed, when this does not hold, we divide $f$ and $g$ by $x^{\text{ord}_{0}{g}}$. This does not change the coefficient sizes and sparsities of $f$ and $g$, and it only reduces their degrees.
4.1 Division by Binomial
In this section, we prove Theorem 1.12. We bound the sparsity of $h$ in terms of $f$ and $g$ when $g$ is a cyclotomic-free binomial. We first deal with the case when $g$ is a linear function. The following lemma is a special case of a result of Lenstra [Len99] that plays an important role in our proofs.
Lemma 4.1 (Proposition 2.3 of [Len99] for the linear case).
Let $f\in\mathbb{Z}[x]$ such that $f=f_{0}+x^{d}f_{1}$. Suppose that
$d-\deg{f_{0}}\geq\frac{1}{\log{2}}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).$
Then, every zero of $f$ in $\mathbb{Q}$ that is not $1$, $-1$ nor 00 is a common zero of both $f_{0}$ and $f_{1}$.
The next argument is similar to the one in [GRT12, Lemma 4.4].
Lemma 4.2.
Let $f,\ell,h\in\mathbb{Z}[x]$ such that $f=\ell h$ and $\ell$ is linear. If $1$ and $-1$ are not roots of $\ell$, then the sparsity of $h$ is at most
$\frac{1}{\log{2}}\left\|{f}\right\|_{0}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).$

### v4

@@ 3 A Bound on the $\ell_{2}$-norm of the Factor Polynomial
We shall need the following definition in order to state our result.
Definition 3.7.
For a polynomial $g(x)=\sum_{i=1}^{s}c_{i}x^{n_{i}}$ with $n_{1}<n_{2}<\ldots<n_{s}$ and $\prod_{i=1}^{s}c_{i}\neq 0$ we define $d(g):=\prod_{i=1}^{s-1}(n_{s}-n_{i})$. When $s=1$ we set $d(g)=1$.
In the next claims we shall use the notation introduce in Section 2.1: $g_{0}=g/x^{\text{ord}_{0}{g}}$ and $g_{\text{rev}}=x^{\deg{g}}\cdot g(1/x)$.
Claim 3.8.
Let $g$ as in 3.7. We have $d(g)=d(g_{0})=\deg{g_{0}}\cdot d((g_{0})^{\prime})$.
Proof.
The claim follow by easy calculations:
$d(g_{0})=\prod_{i=1}^{s-1}\left((n_{s}-n_{1})-(n_{i}-n_{1})\right)=\prod_{i=1}^{s-1}(n_{s}-n_{i})=d(g).$
Next, observe that
$(g_{0})^{\prime}=(\sum_{i=1}^{s}c_{i}x^{n_{i}-n_{1}})^{\prime}=\sum_{i=2}^{s}(n_{i}-n_{1})c_{i}x^{n_{i}-n_{1}-1}.$
Hence,
$\deg{g_{0}}\cdot d((g_{0})^{\prime})$ $=(n_{s}-n_{1})\cdot\prod_{i=2}^{s-1}\left((n_{s}-n_{1}-1)-(n_{i}-n_{1}-1)\right)$
$=\prod_{i=1}^{s-1}(n_{s}-n_{i})=d(g).\qed$
Claim 3.9.
Let $g$ as in 3.7. Then, $d(g_{\text{rev}})=\prod_{i=2}^{s}(n_{i}-n_{1})$.
Proof.
The case $s=1$ is clear. Let $m_{s-i+1}=n_{s}-n_{i}$. Clearly, $0=m_{1}<m_{2}<\ldots<m_{s}=n_{s}-n_{1}$. Then, $g_{\text{rev}}=\sum_{i=1}^{s}c_{i}x^{n_{s}-n_{i}}=\sum_{i=1}^{s}c_{s-i+1}x^{m_{i}}$ and we get that
$d(g_{\text{rev}})=\prod_{i=1}^{s-1}(m_{s}-m_{i})=\prod_{i=1}^{s-1}\left((n_{s}-n_{1})-(n_{s}-n_{s-i+1})\right)=\prod_{i=2}^{s}(n_{i}-n_{1}).\qed$
Corollary 3.10.
Let $g$ as in 3.7. Then,
$\max\{d(g),d(g_{\text{rev}})\}\geq(n_{s}-n_{1}-1)^{s/2}.$
Proof.
Note that
$d(g)d(g_{\text{rev}})=\prod_{i=1}^{s-1}(n_{s}-n_{i})\cdot\prod_{i=2}^{s}(n_{i}-n_{1})=(n_{s}-n_{1})^{2}\cdot\prod_{i=2}^{s-1}\left((n_{s}-n_{i})(n_{i}-n_{1})\right).$ (3.11)
Since $(n_{s}-z)(z-n_{1})\geq n_{s}-n_{1}-1$ for any integer $n_{1}<z<n_{s}$, we can conclude that,
$d(g)d(g_{rev})\geq(n_{s}-n_{1})^{2}(n_{s}-n_{1}-1)^{s-2}\geq(n_{s}-n_{1}-1)^{s}.\qed$
We can improve this estimate by a slightly more careful analysis.
Claim 3.12.
Let $g$ as in 3.7. If $n_{s}-n_{1}>1$ then
$\max\{d(g),d(g_{\text{rev}})\}\geq\left(\frac{s}{2e}\right)^{\frac{s}{2}-1}\cdot\left(\frac{n_{s}-n_{1}}{e}\right)^{s/2}.$
Proof.
As in the proof of 3.10 we consider (3.11). Since the minimum of the expression $(n_{s}-z)(z-n_{1})$ for an integer $n_{1}<n_{1}+a\leq z\leq n_{s}-b<n_{s}$ is attained at $z=n_{1}+a$ if $a\leq b$ and at $n_{s}-b$ otherwise, and given that $n_{1},\ldots,n_{s}$ are distinct, we conclude by induction that,
$\max\{d(g),d(g_{\text{rev}})\}\geq\sqrt{d(g)\cdot d(g_{\text{rev}})}$
$\geq(n_{s}-n_{1})\cdot\sqrt{\prod_{i=2}^{\lfloor\frac{s}{2}\rfloor}(n_{s}-n_{1}-i+1)(i-1)\cdot\prod_{i=1}^{\lceil\frac{s}{2}\rceil-1}(n_{s}-n_{1}-i)i}$
$\geq\left(\left\lfloor\frac{s}{2}\right\rfloor-1\right)!\cdot\left(\frac{(n_{s}-n_{1})!}{(n_{s}-n_{1}-\left\lfloor s/2\right\rfloor)!}\right)\cdot\left(\left(n_{s}-n_{1}-\frac{s-1}{2}\right)\left(\frac{s-1}{2}\right)\right)^{\frac{\lceil s/2\rceil-\left\lfloor s/2\right\rfloor}{2}}$
$\geq^{(*)}\left(\frac{s-3}{2e}\right)^{\frac{s}{2}-1}\cdot\left(\left(\frac{n_{s}-n_{1}}{e}\right)^{n_{s}-n_{1}}\cdot\left(\frac{e}{n_{s}-n_{1}-\lfloor s/2\rfloor}\right)^{n_{s}-n_{1}-\lfloor s/2\rfloor}\right)\cdot\left(n_{s}-n_{1}-\frac{s-1}{2}\right)^{\frac{\lceil s/2\rceil-\left\lfloor s/2\right\rfloor}{2}}$
$\geq\left(\frac{s-3}{2e}\right)^{\frac{s}{2}-1}\cdot\left(\frac{n_{s}-n_{1}}{e}\right)^{\lfloor s/2\rfloor}\cdot\left(1+\frac{\lfloor s/2\rfloor}{n_{s}-n_{1}-\lfloor s/2\rfloor}\right)^{n_{s}-n_{1}-\lfloor s/2\rfloor}\cdot\left(n_{s}-n_{1}-\frac{s-1}{2}\right)^{\frac{\lceil s/2\rceil-\left\lfloor s/2\right\rfloor}{2}}$
$\geq\left(\frac{s-3}{e}\right)^{\frac{s}{2}-1}\cdot\left(\frac{n_{s}-n_{1}}{e}\right)^{\lfloor s/2\rfloor}\cdot\left(n_{s}-n_{1}-\frac{s-1}{2}\right)^{\frac{\lceil s/2\rceil-\left\lfloor s/2\right\rfloor}{2}}$
$\geq\left(\frac{s-3}{e}\right)^{\frac{s}{2}-1}\cdot\left(\frac{n_{s}-n_{1}}{e}\right)^{s/2},$
where to deduce inequality ()(*)( ∗ ) we applied Robbins’ lower bound for the factorial function [Rob55] and some simple calculations. We note that for $s<6$ and $n_{s}-n_{1}>1$ 3.10 implies that
$\max\{d(g),d(g_{\text{rev}})\}\geq(n_{s}-n_{1}-1)^{s/2}\geq\left(\frac{n_{s}-n_{1}}{e}\right)^{s/2}.$
Consequently, if $n_{s}-n_{1}>1$ then
$\max\{d(g),d(g_{\text{rev}})\}\geq\left(\frac{s}{2e}\right)^{\frac{s}{2}-1}\cdot\left(\frac{n_{s}-n_{1}}{e}\right)^{s/2}.\qed$
@@
Let $g\in\mathbb{C}[x]$ be a monic polynomial. Then, there exists a set $B(g)\subseteq[0,2\pi)$, satisfying $|B(g)|\leq 12(\left\|{g}\right\|_{0}-1)\cdot(\deg{g}-\text{ord}_{0}{g})$, such that for every $\delta>0$, and every $\alpha\in[0,2\pi)$ which is $(\left\|{g}\right\|_{0}-1)\delta$-far from $B(g)$, it holds that $\left|g(e^{i\alpha})\right|\geq d(g)\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}$.
@@
We prove the lemma by induction on $s=\left\|{g}\right\|_{0}$. The claim holds for $s=1$ since clearly $|g(e^{i\alpha})|=1$ for any $\alpha$, when $g(x)=x^{m}$.
For the inductive step, suppose that the claim holds for $s$, and let $g\in\mathbb{C}[x]$ be of sparsity $s+1$. Since $\left|(g/x^{\text{ord}_{0}{g}})(\omega)\right|=\left|g(\omega)\right|$ for every $\omega$ on the unit circle we can consider $g_{0}=(g/x^{\text{ord}_{0}{g}})$ instead of $g$, which is of degree $\deg{g_{0}}=\deg{g}-\text{ord}_{0}{g}$.
Let $f:[0,2\pi)\to\mathbb{C}$ be defined as $f(x)=g_{0}(e^{ix})$. Denote with
$\hat{g}(x)=\frac{g_{0}^{\prime}(x)}{\deg{g_{0}}}\,,$
the monic polynomial which is a scalar multiple of $g_{0}^{\prime}$. By 3.8, $d(\hat{g})=d(g_{0}^{\prime})$. Let
@@
$B(g)$ $=B(\hat{g})\cup\left\{\beta\in[0,2\pi):h(\beta)=0\text{ for a non-zero }h\in H\right\}.$
By 2.2, each non-zero function in $H$ has at most $2\deg{g_{0}}$ zeros in $[0,2\pi)$. Since $\hat{g}(x)$ is monic and has sparsity $s$, we conclude by the inductive assumption that
$\left|B(g)\right|\leq 12(s-1)\deg{\hat{g}}+12\deg{g_{0}}<12s\deg{g_{0}}.$
Let $I=[a,b]\subseteq[0,2\pi)$ be an interval that is $(s-1)\delta$-far from $B(g)$. By the definition of $B(g)$, each functions in the set $H$, as defined in Equation (3.14), does not change its sign within $I$, since any sign change would have to go through a zero of the function. As $f^{\prime}(x)=ie^{ix}\cdot{g_{0}^{\prime}}(e^{ix})$, 3.6 implies that for every $\beta\in[a+\delta,b-\delta]$ it holds that
$\left|g(e^{i\beta})\right|$ $=\left|g_{0}(e^{i\beta})\right|=\left|f(\beta)\right|\geq\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|f^{\prime}(\alpha)\right|$ (3.15)
$=\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|g_{0}^{\prime}(e^{i\alpha})\right|=\frac{\delta\cdot\deg{g_{0}}}{\sqrt{2}}\min_{\alpha\in I}\left|\hat{g}(e^{i\alpha})\right|.$
By definition, $B(\hat{g})\subseteq B(g)$ and as each $\beta\in I$ is at least $(s-1)\delta$-far from $B(g)$, we get from the induction hypothesis applied to $\hat{g}$, Equation (3.15) and 3.8
$\left|g(e^{i\beta})\right|$ $\geq\frac{\delta\cdot\deg(g_{0})}{\sqrt{2}}\min_{\alpha\in I}\left|\hat{g}(e^{i\alpha})\right|$
$\geq\frac{\delta\cdot\deg(g_{0})}{\sqrt{2}}\cdot d(\hat{g})\left(\frac{\delta}{\sqrt{2}}\right)^{s-1}$
$=\frac{\delta\cdot\deg(g_{0})}{\sqrt{2}}\cdot d({g_{0}^{\prime}})\left(\frac{\delta}{\sqrt{2}}\right)^{s-1}$
$=d(g)\left(\frac{\delta}{\sqrt{2}}\right)^{s}.$
In other words, every $\beta$ that is $s\delta$ far from the set $B(g)$ satisfies $\left|g(e^{i\beta})\right|\geq d(g)\left(\frac{\delta}{\sqrt{2}}\right)^{s}$, which proves the step of the induction. ∎
Corollary 3.16.
Let $g\in\mathbb{C}[x]$ be a polynomial. Then, there exists a set $B(g)\subseteq[0,2\pi)$, of size $|B(g)|\leq 12(\left\|{g}\right\|_{0}-1)(\deg{g}-\text{ord}_{0}{g})$, such that for every $\delta>0$ and every $\alpha\in[0,2\pi)$ which is $(\left\|{g}\right\|_{0}-1)\delta$-far from $B(g)$, it holds that
$\left|g(e^{i\alpha})\right|\geq\max\{\left|\text{LC}({g})\right|\cdot d({g}),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}\cdot\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}.$
Proof.
Let $g_{\text{rev}}=x^{\deg{g}}g(1/x)$. Since for every $\alpha$, $\left|g(e^{i\alpha})\right|=\left|g_{\text{rev}}(e^{-i\alpha})\right|$, we can consider either $g$ or $g_{\text{rev}}$ in order to prove the claim (replacing $B(g)$ with $B(g_{\text{rev}})$ if required). Note that $g/\text{LC}(g)$ and $g_{\text{rev}}/\text{LC}(g_{\text{rev}})=g_{\text{rev}}/\text{TC}(g)$ are monic polynomials. From 3.13 we have (where $\alpha$ is $\delta$-far from either $B(g)$ or $B(g_{\text{rev}})$, respectively)
$\left|\frac{g(e^{i\alpha})}{\text{LC}(g)}\right|\geq d(g)\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}$ $\Rightarrow\left|g(e^{i\alpha})\right|\geq\left|\text{LC}(g)\right|\cdot d(g)\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}$
$\left|\frac{g_{\text{rev}}(e^{i\alpha})}{\text{LC}(g_{\text{rev}})}\right|\geq d(g_{\text{rev}})\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}$ $\Rightarrow\left|{g}(e^{i\alpha})\right|\geq\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}.\qed$
@@
Let $g\in\mathbb{C}[x]$. Let $p_{\min{}}\in\mathbb{N}$ and $L(n)=n\log{n}$. Set $p_{\max{}}=\lceil 2L(p_{\min{}}+12\left\|{g}\right\|_{0}(\deg{g}-\text{ord}_{0}{g}))\rceil$. Then, there exists a prime $p\in(p_{\min{}},p_{\max{}}]$ such that
$\left|g(\omega)\right|\geq\max\{\left|\text{LC}({g})\right|\cdot d({g}),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}\cdot\left(\frac{\pi}{\sqrt{2}\cdot\left\|{g}\right\|_{0}\cdot p_{\max{}}^{2}}\right)^{\left\|{g}\right\|_{0}-1},$
@@
If $\left\|{g}\right\|_{0}=1$, then the statement is trivial since $\left|g(\omega)\right|=\left|\text{LC}(g)\right|=\left|\text{TC}(g)\right|$ for every root of unity $\omega$. Hence, we can assume that $\left\|{g}\right\|_{0}\geq 2$ and $\deg{g}\geq 1$.
Let $t=p_{\min{}}+12\left\|{g}\right\|_{0}(\deg{g}-\text{ord}_{0}{g})$, and let $\pi(n)$ be the prime counting function. Our assumption implies that $t\geq 24$. By 2.4 we conclude that
$\pi(p_{\max{}})$ $=\pi(\lceil 2t\log{t}\rceil)$
@@
Let $B(g)$ be as in 3.13 and
$P=\left\{p_{\min{}}<p\leq p_{\max{}}:p\text{ is prime}\right\}.$
We bound the size of $P$ by counting the primes in the interval $(p_{\min{}},p_{\max{}}]$:
$\left|P\right|\geq\pi(p_{\max{}})-p_{\min{}}>t-p_{\min{}}=12\left\|{g}\right\|_{0}(\deg{g}-\text{ord}_{0}{g})\geq\left|B(g)\right|.$
By 3.2, there exists $p\in P$ such that for any integer $0<a<p$ and any $\alpha\in B(g)$
$\left|a/p-\alpha/2\pi\right|\geq\frac{1}{2p_{\max{}}^{2}}.$
@@
$\left|2\pi\frac{a}{p}-\alpha\right|\geq\frac{\pi}{p_{\max{}}^{2}}.$
In particular, every non-trivial $p$th root of unity is $\frac{\pi}{p_{\max{}}^{2}}$-far from $B(g)$. Finally, by applying 3.16 with $\delta=\frac{\pi}{\|g\|_{0}\cdot p_{\max{}}^{2}}$, we conclude that for any $p$th root of unity $\omega\neq 1$,
$\left|g(\omega)\right|\geq\max\{\left|\text{LC}({g})\right|\cdot d({g}),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}\cdot\left(\frac{\pi}{\sqrt{2}\cdot\left\|{g}\right\|_{0}\cdot p_{\max{}}^{2}}\right)^{\left\|{g}\right\|_{0}-1}.\qed$
@@
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$. Denote $L(n)=n\log{n}$. Then, for $\text{M}:=\max\{\left|\text{LC}({g})\right|\cdot d({g}),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}$
$\left\|{h}\right\|_{2}\leq\frac{\sqrt{2}\left\|{f}\right\|_{1}}{\text{M}{}}\left({2\left\|{g}\right\|_{0}\cdot L\left(12\left\|{g}\right\|_{0}(\deg{g}-\text{ord}_{0}{g})+2\deg{f}\right)^{2}}\right)^{\left\|{g}\right\|_{0}-1}.$
@@
Let $p_{\min{}}=2\deg{f}$ and $p_{\max{}}$ as in 3.17. From 3.17 we conclude that there exists a prime $p\in(p_{\min{}},p_{\max{}}]$ such that
$\left|g(\omega)\right|\geq\text{M}\cdot\left(\frac{\pi}{\sqrt{2}\cdot\left\|{g}\right\|_{0}\cdot(\lceil 2L(12\left\|{g}\right\|_{0}(\deg{g}-\text{ord}_{0}{g})+2\deg{f})\rceil)^{2}}\right)^{\left\|{g}\right\|_{0}-1}$
@@
To conclude the proof we prove a lower bound on $\text{M}=\max\{\left|\text{LC}({g})\right|\cdot d({g}),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}$.
Lemma 3.19.
Let $g\in\mathbb{C}[x]$. Let $\text{M}=\max\{\left|\text{LC}({g})\right|\cdot d({g}),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}$. Then,
$\text{M}\geq\max\left\{\begin{array}[]{c}\min\{\left|\text{LC}(g)\right|,\left|\text{TC}(g)\right|\}\cdot\left(\frac{\|g\|_{0}}{2e}\right)^{\frac{\|g\|_{0}}{2}-1}\cdot\left(\frac{\deg{g}-\text{ord}_{0}{g}}{e}\right)^{\frac{\left\|{g}\right\|_{0}}{2}}\\ \max\{\left|\text{LC}({g})\right|,\left|\text{TC}({g})\right|\}\cdot(\deg{g}-\text{ord}_{0}{g})\end{array}\right\}.$
Proof.
Note that $d(g),d(g_{\text{rev}})\geq(\deg{g}-\text{ord}_{0}{g})$ and therefore
$\text{M}\geq\max\{\left|\text{LC}({g})\right|,\left|\text{TC}({g})\right|\}\cdot(\deg{g}-\text{ord}_{0}{g}).$
In addition, since $\max\{\left|\text{LC}(g)\right|\cdot d(g),\left|\text{TC}({g})\right|\cdot d(g_{\text{rev}})\}\geq\min\{\left|\text{LC}(g)\right|,\left|\text{TC}({g})\right|\}\cdot\max\{d(g),d(g_{\text{rev}})\}$, 3.12 implies that
$\text{M}\geq\min\{\left|\text{LC}(g)\right|,\left|\text{TC}(g)\right|\}\cdot\left(\frac{\|g\|_{0}}{2e}\right)^{\frac{\|g\|_{0}}{2}-1}\cdot\left(\frac{\deg{g}-\text{ord}_{0}{g}}{e}\right)^{\frac{\left\|{g}\right\|_{0}}{2}}.\qed$
Theorem 1.8 immediately follows by combining the estimate from 3.19 with Theorem 3.18.
@@
Let $d_{1},d_{2}\in\mathbb{N}$ and $a\in\mathbb{C}$ with $\left|a\right|\geq 1$. We conclude the claim by noting that
$x^{d_{1}}\bmod(x^{d_{2}+1}-ax^{d_{2}})=a^{d_{1}-d_{2}}x^{d_{2}}.\qed$
Corollary 3.21.
In the setting above, if either $g\in\mathbb{C}[x]$ is monic or $g\in\mathbb{Z}[x]$, and in both cases $g(0)\neq 0$, then
$\left\|{h}\right\|_{2}\leq{\sqrt{\frac{\left\|{g}\right\|_{0}}{\deg g}}\left\|{f}\right\|_{1}}\left(\frac{2\left\|{g}\right\|_{0}\cdot L\left(12\left\|{g}\right\|_{0}\deg{g}+2\deg{f}\right)^{2}}{\sqrt{\left(\frac{\left\|{g}\right\|_{0}}{2e}\right)\cdot\left(\frac{\deg{g}}{e}\right)}}\right)^{\left\|{g}\right\|_{0}-1}.$